# Exploring Variational Graph Autoencoders for Distribution Grid Data Generation

### v1

@@ Abstract
To address the lack of public power system data for machine learning research in energy networks, we investigate the use of variational graph autoencoders (VGAEs) for synthetic distribution grid generation. Using two open-source datasets — ENGAGE and DINGO — we evaluate four decoder variants and compare generated networks against the original grids using structural and spectral metrics. Results indicate that simple decoders fail to capture realistic topologies, while GCN-based approaches achieve strong fidelity on ENGAGE but struggle on the more complex DINGO dataset, producing artifacts such as disconnected components and repeated motifs. These findings highlight both the promise and limitations of VGAEs for grid synthesis, underscoring the need for more expressive generative models and robust evaluation. We release our models and analysis as open source to support benchmarking and accelerate progress in ML-driven power system research.
@@ 1 Introduction
The global transition towards decarbonized energy systems is driving a rapid increase in distributed energy resources (DERs) such as photovoltaics, batteries, and electric vehicles. These DERs threaten distribution grid voltage stability and require networks that are flexible, resilient, and able to accommodate highly variable generation and demand patterns. Machine learning is promising in addressing these issues, but research is constrained by a lack of publicly available grid data, since grid operators are reluctant to share detailed information on feeder topologies and device locations due to security and privacy concerns Mohammadi and Saleh (2021). Recent works have highlighted the importance of training data for robust grid models, emphasizing its impact on model generalization Okoyomon and Goebel (2025) and feasibility region coverage Joswig-Jones et al. (2022). This lack of data availability creates a bottleneck for developing and validating new algorithms for reliable and efficient grid operation under increasing renewable penetration.
To overcome this barrier, synthetic grid generation has emerged as a promising avenue. Researchers have proposed methods to construct artificial test systems that preserve the key structural and operational characteristics of real grids. At present, the most widely used approaches for this data generation rely on statistical models and heuristic algorithms Mohammadi and Saleh (2021). Statistical models approximate real-world feeders by sampling from probability distributions of observed grid properties, while heuristic algorithms construct networks by solving optimization problems subject to structural and operational constraints. In practice, both methods induce simplified assumptions that restrict their ability to generate data that captures the diversity, dynamics, and nontrivial relationships that exist in real power systems.

### v2

@@ Abstract
To address the lack of publicly available power system data for machine learning research in energy networks, we investigate the use of variational graph autoencoders (VGAEs) for generating synthetic distribution grids. Access to detailed feeder models is constrained by privacy and security considerations; yet such data is essential for advancing machine learning approaches that enable reliable and flexible operation of distribution grids with high shares of renewable generation. Using two open-source datasets—ENGAGE and DINGO—we evaluate four decoder variants and compare the generated networks against real grids using structural and spectral metrics. Results show that simple decoders fail to capture realistic topologies, while GCN-based methods achieve strong fidelity on ENGAGE but struggle with the more complex DINGO dataset, producing artifacts such as disconnected components and repeated motifs. These findings highlight both the potential and limitations of VGAEs for grid synthesis, underscoring the need for more expressive generative models and robust evaluation. We release our models and analysis as open source to support benchmarking and accelerate progress in ML-driven power system research.
@@ 1 Introduction
The global transition toward decarbonized energy systems is accelerating as societies work to mitigate climate change. This shift is driving a rapid rise in distributed energy resources (DERs) such as photovoltaics, batteries, and electric vehicles, which are essential for reducing carbon emissions. Their inherently variable and decentralized nature, however, introduces new operational challenges: high renewable penetration can threaten voltage stability and requires distribution networks that are flexible, resilient, and capable of managing rapid fluctuations in supply and demand. Advanced data-driven tools, including machine learning, show promise in supporting reliable, low-carbon grid operation; however, research is constrained by a lack of publicly available distribution grid data. Utilities are reluctant to share feeder topologies and device information due to security and privacy concerns Mohammadi and Saleh (2021). Recent studies emphasize that the quality and coverage of training data directly affect model generalization Okoyomon and Goebel (2025) and feasibility exploration Joswig-Jones et al. (2022), making this data scarcity a central barrier to developing climate-aligned grid planning and control algorithms.
To address this limitation, synthetic grid generation has emerged as a compelling solution. Rather than relying on sensitive real-world data, researchers construct artificial test systems that emulate key structural and operational properties of distribution networks. Currently, the most widely used approaches for generating this data rely on statistical models and heuristic algorithms Mohammadi and Saleh (2021). Statistical models approximate real-world feeders by sampling from probability distributions of observed grid properties, while heuristic algorithms construct networks by solving optimization problems subject to structural and operational constraints. In practice, both methods induce simplified assumptions that restrict their ability to generate data that captures the diversity, dynamics, and nontrivial relationships that exist in real power systems.